OpenStudy (amtran_bus):

If a sweater sells for $48 after a 25 % mark down, what is the original price?

OpenStudy (mathmate):

Use proportions. Example: If a shirt sells for $30 after a 20 % mark down, what is the original price? 20% discount means the price is (100-20)=80% of the original price. So using proportions, 80% : $30 100% : X Now cross multiply, i.e. (and replace : by =) 80%*X = $30*100% divide both sides by 80% X=$30*(100%/80%)=$30*100/80=$37.50 (use your calculator if necessary) Proceed the same way with your question and post your answer for a check if you wish.

OpenStudy (amtran_bus):

so 64?